### v1

# RES-Q: Evaluating Code-Editing Large Language Model Systems at the Repository Scale

Abstract
The instruction-following ability of Large Language Models (LLMs) has cultivated a class of LLM-based systems capable of approaching complex tasks such as making edits to large code repositories. Due to the high sensitivity and unpredictability of LLM behavior in response to changes in prompting, robust evaluation tools are needed to drive future iteration of these systems. We propose RES-Q, a natural language instruction-based benchmark for evaluating Repository Editing Systems, which consists of 100 repository editing tasks derived from real GitHub commits. Given an edit instruction and a code repository, RES-Q evaluates an LLM system’s ability to gather information and construct an edit that satisfies the criteria set by the instruction. We argue that evaluating LLMs in this way addresses issues with traditional benchmarks and provides a more holistic assessment of a model’s abilities. We evaluate various state-of-the-art LLMs as language agents in a repository-editing system built on Qurrent OS, our language agent development software. Despite their 1% pass@1 performance difference on HumanEval, we find Claude Sonnet 3.5 outperforms GPT-4o by 12% pass@1 on RES-Q, indicating RES-Q’s capacity to differentiate model capability as traditional benchmarks approach saturation. We further investigate token efficiency, performance relationships with existing benchmarks, and interesting disparities between closed and open-source LLMs. Code and dataset are available at this https URL.
1 Introduction
1.1 Automating Software Development with LLMs
The recent increase in the versatile capability of Large Language Models (LLMs) has given rise to a class of systems referred to as LLM-based agents, or language agents for short (Xi et al., 2023). In general, a language agent is LLM augmented by planning, memory, and tool-use capabilities (Weng, 2023). Under this umbrella, there exists a set of language agent systems that seek to automate software development tasks. Osika (2023) employs a user feedback loop to clarify task details. Significant Gravitas (2024) decomposes its task into a tree of subtasks. Hong et al. (2023) coordinates multiple language agents working as different roles in a software assembly line. Each with a unique strategy, the ability of these systems to develop simple web apps has been an impressive enough spectacle to raise questions about the future role of humans in software development.
However, the majority of software development work involves extending, editing, and maintaining existing large codebases rather than creating new ones from scratch. In these more realistic settings, the challenge of managing and reasoning across large contexts has been prohibitive to similarly sensational performance. Nonetheless, recent efforts such as CodeR (Chen et al., 2024) and Aider (Gauthier, 2023) have shown the potential for near-term improvement in these areas with LLM-based systems. As models and systems become more capable, it is crucial to have robust evaluation frameworks that can accurately assess their performance on real-world software development tasks.
1.2 Evaluation of LLMs
Historically, the performance of an LLM has been characterized by its accuracy on a set of annotated NLP benchmarks such as BIG-Bench Hard (Suzgun et al., 2022) and MMLU (Hendrycks et al., 2021) which test knowledge and problem-solving capabilities. However, these benchmarks are primarily designed to evaluate LLMs in zero-shot or few-shot settings (Hendrycks et al., 2021), measuring only the knowledge acquired by the model during pre-training. Additionally, Ott et al. (2022) shows that the majority of NLP benchmarks quickly tend towards saturation, and Sainz et al. (2023) demonstrates that data contamination can compromise traditional annotated benchmarks, leading to an overestimation of model performance.
Evaluating LLMs based on their performance as a language agent addresses the identified shortcomings of current annotated NLP benchmarks. Specifically, the utility of a particular LLM as a language agent depends on its ability to precisely follow instructions, understand roles, and make informed, sequential decisions, thereby providing a more holistic evaluation of the model. Additionally, complex tasks undertaken by language agents inherently present higher difficulty levels, reducing the risk of benchmark saturation observed in traditional evaluations. Furthermore, the novelty and diversity of language agent trajectories significantly mitigate the risk of data contamination, as each trajectory is influenced by unique interactions and evolving contexts.
1.3 Contributions
In this paper, we introduce RES-Q, a hand-crafted dataset consisting of 100 natural language code repository edit instructions paired with corresponding test suites. RES-Q’s instructions are derived from actual commits made to GitHub repositories, ensuring their relevance and realism. Concretely, RES-Q serves as a robust evaluation tool for assessing the performance of LLM-based systems on real-world software development tasks, particularly those involving the extension and modification of large code repositories. More generally, RES-Q provides an opportunity to evaluate the objective, binary success of an LLM working as a language agent on a complex task, thereby offering a more comprehensive indication of its capabilities compared to traditional benchmarks. To this end, we conduct an assessment of state-of-the-art models within a code-repository editing system built on Qurrent OS, our proprietary LLM agent development software, and compare the results with traditional benchmarks. To complement our dataset, we also release a submission environment to evaluate the correctness of repository changes for a given task, promoting broader community engagement.
To summarize, we make the following contributions:
  1. 1.
    Introduction of RES-Q Dataset: We introduce RES-Q, an instruction-based dataset of codebase edits derived from actual GitHub commits, designed to evaluate the performance of LLM-based systems on real-world software development tasks (§ 3).
  2. 2.
    Evaluation of SOTA LLMs: We evaluate state-of-the-art LLMs within a codebase-editing system built on Qurrent OS, our LLM agent development software, and investigate correlations with performance on traditional benchmarks (§ 4).
  3. 3.
    Submission Environment: We release a submission environment with a simple interface for evaluating the correctness of codebase changes for a given RES-Q task, facilitating the evaluation of other systems and encouraging broader community engagement.
2 Related Work
Code Editing Benchmarks. SWE-Bench (Jimenez et al., 2024) is a Python codebase editing benchmark constructed automatically from GitHub Issues and their accompanying Pull Requests. Designed to closely emulate real-world software development tasks, SWE-Bench uses raw GitHub Issue threads as task instructions. In contrast, RES-Q’s instructions consist of intentionally vague, mostly single-sentence imperative instructions, similar to the lazy edit instructions described in Cassano et al. (2024). While GitHub Issues are indeed a realistic artifact of software development, engineers initially approaching a task will not often have the prescriptiveness provided by GitHub Issues. RES-Q’s instructions capture this reality by testing the ability to reason with ambiguity present in task definitions. Additionally, by avoiding the direct use of GitHub data in task instructions, RES-Q reduces the risk of data contamination for models trained on GitHub. SWE-Bench filters for task instances with fail-to-pass tests, unit tests introduced in the PR that fail on the repository before the PR and pass after (Jimenez et al., 2024). While this ensures that changes meet a minimum threshold of correctness, it does not guarantee comprehensive verification. Unit tests added during PRs often cover specific edge cases or a limited subset of the new functionality, potentially leaving significant portions of the code unverified. In contrast, RES-Q includes a meticulously handcrafted test suite for each task to ensure exhaustive verification of solutions. Moreover, while SWE-Bench focuses solely on Python codebases, RES-Q includes tasks for both Python and JavaScript codebases, broadening the scope of evaluation.
CanItEdit (Cassano et al., 2024) is a handcrafted dataset of 105 instructional code editing problems, each centered on single Python files. CanItEdit’s lazy instructions model is characterized by its intentional brevity and vagueness, an approach that has been adopted by RES-Q. Both benchmarks use handcrafted test suites to validate task solutions. CanItEdit explicitly provides the file to be edited, isolating the precise code editing task within a single file context. In contrast, RES-Q provides the entire codebase as context, requiring deliberation about which files need to be edited, and thereby presenting a broader software development task for evaluation. In terms of task complexity, CanItEdit tasks involve changing only one file, while RES-Q tasks often require edits across multiple files. Additionally, the edits in RES-Q are generally more extensive and complex compared to those in CanItEdit, as indicated by the higher median Levenshtein distance (see Table 1). This highlights RES-Q’s focus on broader context tasks, aligning more closely with real-world software development challenges.
Overall, RES-Q integrates the commit-based, handcrafted nature of CanItEdit with the real-world task environment of SWE-bench. This combination aims to provide a comprehensive and realistic evaluation framework for LLM-based systems, facilitating their assessment on practical software development tasks and the broader capabilities of the system’s LLM.
3 RES-Q
Figure 1: Distribution of RES-Q tasks across different repositories.
Table 1: Dataset statistics for RES-Q.
RES-Q Dataset Statistics
Total Tasks 100
Repository Q1 $\mid$ Median $\mid$ Q3
Lines of Code (LoC) 4170 $\mid$ 6468 $\mid$ 42793
Files 20 $\mid$ 55 $\mid$ 395
Ground Truth Patch Q1 $\mid$ Median $\mid$ Q3
Levenshtein Distance 119.5 $\mid$ 730 $\mid$ 1720
Modified LoC 6 $\mid$ 27.5 $\mid$ 63.7
Modified Files 1 $\mid$ 1 $\mid$ 3
Tokens 224.5 $\mid$ 523 $\mid$ 885.25
Instruction Mean $\pm$ Std. Dev.
Tokens 70.45 $\pm$ 36.68
RES-Q is a codebase editing benchmark based on compact, natural language instructions. Given an edit instruction and a codebase, the task is to make an edit to the codebase that satisfies the criteria set by the instruction. Table 1 presents general dataset statistics for RES-Q.
3.1 Benchmark Construction
To construct RES-Q, we selected 14 GitHub repositories of varying size and popularity. We combed through each repository’s commit history and selected a representative collection of changes whose distribution mirrors real-world software development practices. The distribution of RES-Q tasks across these repositories is shown in Figure 1, and Figure 3 illustrates the distribution of tasks across different classifications. For each selected change, we recorded the base commit (the starting point of the task), as well as its associated solution commit, (the ground truth change). It’s important to note that the solution commit is not necessarily the direct child of the base commit; intermediate commits may exist between them.
We crafted an edit instruction and a Python-based testing environment for each selected change. Each testing environment includes a Python version, a requirements file, and a test suite. As an additional verification step, we used our submission environment to ensure that the test suite fails on the base commit, and passes on a patch generated from the diff between the base commit and solution commit.
We created forks of each repository included in the benchmark to ensure consistency over time.
3.2 Task Formulation
Figure 2: Distribution of RES-Q tasks across different instruction types.
Figure 3: The 8 most frequent primary action verbs with their corresponding 4 most frequent target nouns occurring in RES-Q task instructions.
Prompt. The prompt for each RES-Q task includes:
  1. 1.
    A natural language instruction describing the desired change.
  2. 2.
    The entire codebase checked out at the task’s base commit.
In practice, the RES-Q dataset provides the forked GitHub repository URL and base commit hash as a reference to the codebase.
Submission. A RES-Q task submission consists of an edit to the codebase which makes the change described in the instruction. The submission environment, described in § 3.3, represents this edit as a unified diff patch file.
Evaluation Metric. Tasks are evaluated based on binary success. A task is considered successful if the provided patch applies successfully to the repository, and the resulting repository passes the task’s test suite. The primary metric for RES-Q is pass@k which measures the likelihood of producing at least one successful completion per task, given a fixed number k attempts. In the case where each task is attempted once, pass@1 represents the proportion of successful tasks to total tasks.
An example RES-Q task is provided in § A.1.
3.3 Submission Environment
We implement a submission environment to facilitate the easy evaluation of solutions to RES-Q tasks generated by other LLM systems. The environment accepts a submission, consisting of a task identifier and unified diff patch file, and produces a boolean indicating the success of the submission, a feedback message (one of PASS, FAIL, PATCH FAILED, TIMED OUT), and feedback from the test suite. The environment is implemented asynchronously, enabling concurrent evaluation of multiple submissions while maintaining complete isolation of each one. This design ensures both scalability and reliability in the evaluation process. See § A.3 for a diagram describing the high-level architecture of the submission environment.
4 Evaluation
(a) RES-Q vs HumanEval
(b) RES-Q vs MMLU
Figure 4: Model performances on RES-Q vs traditional LLM benchmarks. Note: The x-axes (HumanEval and MMLU scores) and y-axes (RES-Q scores) use different scales to better highlight differences between models.
In this section, we evaluate the performance of a set of state-of-the-art open and closed source LLMs on the RES-Q benchmark.
4.1 Experimental Setup
To produce submissions for RES-Q tasks, we implemented an autonomous, multi-agent repository-editing system on top of Qurrent OS, our proprietary LLM agent development software. The system integrates code-editing tool use and an agent-based code retrieval system. Given an edit instruction and a Git repository, the system makes edits directly to the repository. We use the patch generated from the diff between the base commit and the edits made by the system as our submission to the RES-Q submission environment.
We selected various state-of-the-art closed and open source LLMs for evaluation. We attempt all 100 RES-Q tasks once with each selected LLM acting in the repository-editing system and report their pass@1 scores.
Due to the wide range of context sizes in the models selected for evaluation (see Table 2), we evaluated in two settings: (1) Unconstrained: each language agent is afforded its model’s entire context window. (2) Constrained: each language agent is afforded 8K tokens of its model’s context window. In the first setting, we evaluate the full capacity of each model and its ability to capitalize on its context window. In the second setting, we standardize window size, isolating evaluation of the model’s per-token ability.
4.2 Results
Constraining Context Window. For the selected closed-source models, constraining the context window size to 8K tokens consistently decreased performance on RES-Q. The largest performance decreases by constraining context were observed in the two top-performing models: Claude 3.5 Sonnet (-9%) and GPT-4o (-10%). For the selected open-source models, constraining the context window consistently increased performance. Qualitatively, this indicates that these models are not able to utilize their entire context window as effectively. Relative performance of models remained unchanged across the constrained and unconstrained settings.
Closed vs Open Source LLMs. With the exception of DeepSeek Coder V2 (29% pass@1) which outperformed Gemini Pro 1.5 (27% pass@1) in the constrained setting, closed-source LLMs outperformed open-source LLMs on RES-Q across the board. Most notably, Claude Sonnet 3.5 outperformed all other models including GPT-4o by 12% pass@1 in the unconstrained setting and 15% pass@1 in the constrained setting.
Performance Correlation with Existing Benchmarks. Our analysis identified two statistically significant performance correlations ($p<0.05$) between RES-Q and the selected traditional benchmarks. When considering all selected LLMs, the correlation between RES-Q and MMLU was strong with a coefficient of 0.73 ($p=0.00054$) while the correlation between RES-Q and HumanEval was not significant ($p=0.054$). When considering only closed-source LLMs, the correlation between RES-Q and HumanEval was very strong with a coefficient of 0.86 ($p=0.0016$) while the correlation between RES-Q and MMLU was not significant ($p=0.065$). We did not identify significant performance correlations when considering only open-source models.
Per-Task Token Usage. We found that, despite its superior performance, Claude Sonnet 3.5 had the highest median token usage per RES-Q task in both the constrained and unconstrained settings. Qualitatively, this could indicate that Claude may make mistakes early on in a trajectory, but is able to recover and successfully realign to a task. On the other hand, GPT-4o uses a disproportionately low amount of tokens per task ($\sim$ 50% that of Sonnet), given its performance. This may indicate that it is better at making correct decisions early on in a trajectory.
Figure 5: Median Input and Output Tokens Used Per Task by Model in each of the Constrained (8K window) and Unconstrained Settings.
Model Benchmark Results (pass@1)
Name Context Window HumanEval MMLU RES-Q RES-Q (8K)
Closed Models
Claude 3.5 Sonnet 200K 92.0 88.3 58 51
GPT-4o 128K 91 88.7 46 36
GPT-4T 128K 88.2 86.5 37 34
Claude 3 Opus 200K 84.8 86.8 36 32
GPT-4 8K 82.68 86.4 30 30
Gemini 1.5 Pro 1M 71.9 81.9 30 27
Open Models
DeepSeek Coder V2 Instruct 128K 90.2 79.2 26 29
Llama 3 70b 8K 81.7 82 20 20
Qwen2-72B Instruct 32K 86.0 82.3 16 18
Table 2: Performance of selected LLMs on RES-Q vs traditional NLP benchmarks along with their context window size. Results reported for RES-Q are the product of a language agent system built on Qurrent OS. For details on the API endpoints of closed models, see Appendix A.2.
Appendix A Appendix
A.1 Example RES-Q Task
Instruction Enhance the HTTPClient class in http.py to allow editing various application details via Discord’s API. Implement the edit_application_info method to send a PATCH request to the /applications/@me endpoint. This method should accept reason and payload, filter the payload for specific valid keys like ’custom_install_url’, ’description’, etc., and use self.request to send the PATCH request with the filtered payload and reason. Ensure the method returns the response from the API call.
Ground Truth Patch diff --git a/discord/http.py b/discord/http.py --- a/discord/http.py +++ b/discord/http.py @@ -2456,6 +2456,22 @@ class HTTPClient: def application_info(self) -> Response[appinfo.AppInfo]: return self.request(Route(’GET’, ’/oauth2/applications/@me’)) + def edit_application_info(self, *, reason: Optional[str], payload: Any) -> Response[appinfo.AppInfo]: + valid_keys = ( + custom_install_url’, + description’, + role_connections_verification_url’, + install_params’, + flags’, + icon’, + cover_image’, + interactions_endpoint_url’, + tags’, + ) + + payload = {k: v for k, v in payload.items() if k in valid_keys} + return self.request(Route(’PATCH’, ’/applications/@me’), json=payload, reason=reason) +
A.2 API Endpoints for Closed Models
Model API Endpoint
GPT-4o gpt-4o
GPT-4T gpt-4-turbo-2024-04-09
GPT-4 gpt-4-0613
Claude 3 Opus claude-3-opus-20240229
Claude 3.5 Sonnet claude-3-5-sonnet-20240620
Gemini 1.5 Pro gemini-1.5-pro-001
Table 3: API endpoints for the closed models used in our evaluations.
A.3 Submission Environment

### v2

@@
 Due to the high sensitivity and unpredictability of LLM behavior in response to changes in prompting, robust evaluation tools are needed to drive future iteration of these systems.
-We propose RES-Q, a natural language instruction-based benchmark for evaluating Repository Editing Systems, which consists of 100 repository editing tasks derived from real GitHub commits.
-Given an edit instruction and a code repository, RES-Q evaluates an LLM system’s ability to gather information and construct an edit that satisfies the criteria set by the instruction.
+We propose RES-Q, a natural language instruction-based benchmark for evaluating Repository Editing Systems, which consists of 100 handcrafted repository editing tasks derived from real GitHub commits.
+Given an edit instruction and a code repository, RES-Q evaluates an LLM system’s ability to interpret the instruction, navigate the repository to gather relevant information, and construct an appropriate edit that satisfies the specified criteria.
 We argue that evaluating LLMs in this way addresses issues with traditional benchmarks and provides a more holistic assessment of a model’s abilities.
@@
 More generally, RES-Q provides an opportunity to evaluate the objective, binary success of an LLM working as a language agent on a complex task, thereby offering a more comprehensive indication of its capabilities compared to traditional benchmarks.
-To this end, we conduct an assessment of state-of-the-art models within a code-repository editing system built on Qurrent OS, our proprietary LLM agent development software, and compare the results with traditional benchmarks.
+To this end, we conduct an assessment of state-of-the-art models within a code-repository editing system built on Qurrent OS, our proprietary language agent development software, and compare the results with traditional benchmarks.
 To complement our dataset, we also release a submission environment to evaluate the correctness of repository changes for a given task, promoting broader community engagement.
@@
   2. 2.
-    Evaluation of SOTA LLMs: We evaluate state-of-the-art LLMs within a codebase-editing system built on Qurrent OS, our LLM agent development software, and investigate correlations with performance on traditional benchmarks (§ 4).
+    Evaluation of SOTA LLMs: We evaluate state-of-the-art LLMs within a codebase-editing system built on Qurrent OS, our language agent development software, and investigate correlations with performance on traditional benchmarks (§ 4).
   3. 3.
-    Submission Environment: We release a submission environment with a simple interface for evaluating the correctness of codebase changes for a given RES-Q task, facilitating the evaluation of other systems and encouraging broader community engagement.
+    Submission Environment: We release a submission environment with a simple interface for evaluating the correctness of codebase changes for a given RES-Q task, facilitating the evaluation of other systems and encouraging broader community engagement (§ 3.3).
 2 Related Work
@@
 For the selected open-source models, constraining the context window consistently increased performance.
-Qualitatively, this indicates that these models are not able to utilize their entire context window as effectively.
+This suggests that these models may not be able to utilize their entire context window as effectively.
 Relative performance of models remained unchanged across the constrained and unconstrained settings.